OpenStudy (anonymous):

Simplify the expression sin2 t csc2 t − cos2 t

OpenStudy (anonymous):

= 2sintcost / 2sintcost - (2cos^2t - 1) = 1 - 2cos^2t + 1 = 2 - 2cos^2t = 2 (1 - cos^2t) = 2 sin^2t